(5x-3)^2/3 =9

the ^2/3 is an fraction exponent**
1
Expert's answer
2012-03-28T15:46:31-0400
(5x-3)^2/3 =9
(5x-3)^2 =9^3
5x-3=+-9^(3/2)
5x=3+-9^(3/2) =3+-27
x=(3+-27)/5
So we get x=(3+27)/5=6 or x=(3-27)/5=-24/5